# In-ice Askaryan Emission from Air Showers: Implications for Radio Neutrino Detectors

### v1

@@ 4.4 Verification through direct signal trajectories
We show the expected detection rates of air-shower cores with direct signal trajectories as a function of cosmic-ray energy in Fig. 11. We note that the rate prediction only depends on the factor. We find that the detection rate at the IceCube-Gen2 detector at the South Pole is significantly larger than for the RNO-G detector at Summit Station in Greenland. This is because the deep component is placed at a depth of at the South Pole and only at Summit Station. In addition, the index-of-refraction profile is different. The transition from snow to deep ice, i.e., the firn, is in the upper at the South Pole, whereas in Greenland, this transition already occurs within the first . The index-of-refraction gradient bends signal trajectories downwards. The RNO-G phased-array antennas are below the firn layer. Hence, the trajectories are maximally curved, which, in turn, minimizes the physical radius between the core impact location and the RNO-G antenna. Therefore, at the South Pole a larger surface area around the deep antennas has direct signal trajectories: For Gen2 at the South Pole, most direct signal trajectories from air-shower cores are observed from a radius of with the distribution extending to over , while for RNO-G at Summit Station Greenland, the distribution peaks at a radius of with the distribution extending to . Because the flux increases linearly with area, the detection rate for Gen2 is larger than for RNO-G. We note that the additional path length to the deeper antenna has negligible impact because the attenuation length close to the surface where the ice is colder is much larger than the path length.
@@
The total number of events per year and station as a function of the Askaryan Loss factor is shown in Fig. 12. For the reference value of , we expect 10 detections per year and station at RNO-G, which makes this a clearly detectable signal even with a partly deployed detector. Fig. 11 shows that the rate is dominated by lower energy cosmic rays between and in contrast to the reflected rates which are dominated by the end of the cosmic-ray spectrum, i.e., energies larger than eV (cf. Fig. 8). Hence, we expect the average -factor to be smaller for the direct-illumination rates, which could lower the rate predictions of Fig. 11. In the future, we plan to build a better model for the in-ice shower core emission through detailed CORSIKA 8 simulations to decrease this uncertainty.
@@
A deep phased array at the South Pole should see ten times more events, i.e., 100 detections per station per year. One station of the existing ARA detector is equipped with a phased array [17], but without having access to details of the ARA signal chain and trigger performance, we can only approximate its performance. The ARA phased array is deployed at a depth of and consists of 7 antennas, which will increase its performance compared to the four-channel phased array at that we simulated in this study. However, the ARA phased-array operates at a much larger bandwidth which decreases its performance, especially to the low-frequency dominated signals of air-shower cores [18]. Furthermore, the ARA trigger rate is lower (around vs. ); thus, the thresholds are higher, and the signal chain is more dispersive, which reduces performance compared to our estimates for Gen2. Overall, we expect the ARA phased-array sensitivity to be similar but slightly smaller compared to Gen2. As the ARA phased array has been running for several years, the experiment should have hundreds of direct detections of air-shower cores in their dataset. So far, no dedicated search for air-shower core signals has been published, but as a side product of the search for UHE neutrinos [51], 46 impulsive events were found originating from the surface region in a half-year data sample [52] of which a subset could originate from air-shower cores which is well compatible with our rate estimate.
Hence, the Askaryan emission from air-shower cores can be probed experimentally using existing instrumentation. We also show the rate as a function of the shower-core energy and the air shower’s zenith angle in Fig. 13. Both of these quantities can be reconstructed from measurements and compared to our theory prediction.
@@ Acknowledgements
We thank the CORSIKA 8 collaboration for their support. We thank Philipp Windischhofer for helpful discussions about Fresnel reflection coefficients. CG and AC are supported by the Swedish Research Council (Vetenskapsrådet) under project no. 2021-05449. CG is supported by the European Union (ERC, NuRadioOpt, 101116890). This work used resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at UPPMAX partially funded by the Swedish Research Council through grant agreements no. 2022-06725 and no. 2018-05973.

### v2

@@ 4.4 Verification through direct signal trajectories
We show the expected detection rates of air-shower cores with direct signal trajectories as a function of cosmic-ray energy in Fig. 11. We note that the rate prediction only depends on the factor. We find that the detection rate at the IceCube-Gen2 detector at the South Pole is significantly larger than for the RNO-G detector at Summit Station in Greenland. This is because the deep component is placed at a depth of at the South Pole and only at Summit Station. In addition, the index-of-refraction profile is different. The transition from snow to deep ice, i.e., the firn, is in the upper at the South Pole, whereas in Greenland, this transition already occurs within the first . The index-of-refraction gradient bends signal trajectories downwards. The RNO-G phased-array antennas are below the firn layer. Hence, the trajectories are maximally curved, which, in turn, minimizes the physical radius between the core impact location and the RNO-G antenna. Therefore, at the South Pole a larger surface area around the deep antennas has direct signal trajectories: For Gen2 at the South Pole, most direct signal trajectories from air-shower cores are observed from a radius of with the distribution extending to over , while for RNO-G at Summit Station Greenland, the distribution peaks at a radius of with the distribution extending to . Because the flux increases linearly with area, the detection rate for Gen2 is larger than for RNO-G. We note that the additional path length to the deeper antenna has negligible impact because the attenuation length is much larger than the path length.
@@
The total number of events per year and station as a function of the Askaryan Loss factor is shown in Fig. 12. For the reference value of , we expect 10 detections per year and station at RNO-G, which makes this a clearly detectable signal even with a partly deployed detector. Fig. 11 shows that the rate is dominated by lower energy cosmic rays between and in contrast to the reflected rates which are dominated by the end of the cosmic-ray spectrum, i.e., energies larger than eV (cf. Fig. 8). Hence, we expect the average -factor to be smaller for the direct-illumination rates, which could lower the rate predictions of Fig. 11. In the future, we plan to build a better model for the in-ice shower core emission through detailed CORSIKA 8 simulations to decrease this uncertainty. We also show the rate as a function of the shower-core energy and the air shower’s zenith angle in Fig. 13. Both of these quantities can be reconstructed from measurements and compared to our theory prediction.
@@
A deep phased array at the South Pole should see ten times more events, i.e., 100 detections per station per year. One station of the existing ARA detector is equipped with a phased array [22], but without having access to details of the ARA signal chain and trigger performance, we can only approximate its performance. The ARA phased array is deployed at a depth of and consists of 7 antennas, which will increase its performance compared to the four-channel phased array at that we simulated in this study. However, the ARA phased-array operates at a much larger bandwidth which decreases its performance, especially to the low-frequency dominated signals of air-shower cores [23]. Furthermore, the ARA trigger rate is lower (around vs. ); thus, the thresholds are higher, and the signal chain is more dispersive, which reduces performance compared to our estimates for Gen2. Overall, we expect the ARA phased-array sensitivity to be similar but slightly smaller compared to Gen2. As the ARA phased array has been running for several years, the experiment could have hundreds of direct detections of air-shower cores in their dataset. So far, no dedicated search for air-shower core signals has been published, but as a side product of the search for UHE neutrinos [57], 46 impulsive events were found originating from the surface region in a half-year data sample [58] of which a subset could originate from air-shower cores which is well compatible with our rate estimate, but a more thorough calculation for the ARA detector response and trigger settings are required to solidify the expected detection rate. Hence, the Askaryan emission from air-shower cores can be probed experimentally using the existing instrumentation or RNO-G or ARA.
@@ Acknowledgements
We thank the CORSIKA 8 collaboration for their support. We thank Philipp Windischhofer for helpful discussions about Fresnel reflection coefficients. CG and AC are supported by the Swedish Research Council (Vetenskapsrådet) under project no. 2021-05449. CG is supported by the European Union (ERC, NuRadioOpt, 101116890). This work used resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at UPPMAX partially funded by the Swedish Research Council through grant agreements no. 2022-06725 and no. 2018-05973. DB is supported by the National Science Foundation’s generous IceCube EPSCoR Initiative grant #2019597.